# Do Current Language Models Support Code Intelligence for R Programming Language?

### v2

@@ 1. Introduction
However, even though the R language has unparalleled importance in related fields, very limited research on code intelligence in R has been shown. A conventional GPT-2 (Radford et al., 2018) model is used for the code completion task in R (Popov et al., 2021). Though a large number of pre-trained models have been successfully developed in recent years and have proven their effectiveness in various software engineering (SE) tasks (Liu et al., 2019; Devlin et al., 2019; Feng et al., 2020; Guo et al., 2020; Wang et al., 2021b), to the best of our knowledge, no model has attempted to extend their effectiveness to the R language. While progress has been made in code intelligence for R (Popov et al., 2021), there exists a gap in the literature, limiting the realization of its full potential.
@@
In this study, we intend to explore to what extent the current techniques are applicable to R. However, as there is no current dataset, we first collect a dataset of high-quality R packages from GitHub repositories (inclusion/exclusion criteria described in Section 3.2.2). Then, we consider four pre-trained models, CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2020), UniXcoder (Guo et al., 2022) and CodeT5 (Wang et al., 2021b) to empirically study their performance for two tasks of code summarization and method name prediction, with various programming styles and settings. The results of the study are compared across models and with other programming languages. We then investigate an R-specific approach based on different programming styles to evaluate its effect on the models’ performance. In this study we also investigate the performance of LLMs for R language on code generation. This paper is the second Stage of our accepted TOSEM Registered Paper [^4] https://dl.acm.org/journal/tosem/registered-papers, in which we are reporting the results along with the discussions.
Our results reveal several key findings regarding the models and their capabilities for R programming language. First, studied models consistently experience varying degrees of performance degradation when processing R code compared to other programming languages. Second, our experiments demonstrate that multi-language fine-tuning does not uniformly improve performance across all PLMs for R-specific tasks, contrary to expectations. Third, we find that different syntax of the R styles in Tidy-verse and Base code significantly impact the performance, which is affecting the results more for code summarization task. The challenges extend beyond syntax to project-specific contexts inherent in R codebases, which poses a substantial challenge when attempting cross-project training. This is further complicated by the observation that combined syntax approaches generally underperform compared to either individual syntax paradigms. Our experiments with LLMs also show the low Pass@K scores for code generation, falling much behind the scores obtained for Python. These insights emphasize the need for more nuanced approaches to improving model performance for R programming, beyond simply increasing training data or applying multi-language fine-tuning strategies.
@@
It is worth noting that though the foundation models (i.e., Large Language Models) in the past few months have found their way into SE, there is rarely a study or application available to use them for R. Additionally, their usage is costly if one wants to call their APIs. Therefore, their usage is not available for all, and fine-tuning them is not possible due to being large, besides their hidden ability for this programming language. LLMs have some shortcomings (Miceli-Barone et al., 2023) and still studying the smaller Code-PLMs can benefit the community. Therefore, we have studied the Code-PLMs and also added results for code generation using LLMs.
Significance: This study provides valuable insights into the effectiveness of using pre-trained code models for R and highlights the need for further research in this field. The results of our work can be used as a reference for future research and provide a foundation for the development of intelligent tools for the R community. Applying software engineering practices such as version control, documentation, testing, and modularization can lead to more reliable, maintainable, and scalable R code (Vidoni, 2021). Additionally, scientific software developers consider functional correctness, performance, portability, and maintainability as the top important characteristics (Johanson and Hasselbring, 2018). This is of specific importance as research shows that many scientists develop/use scientific software and “better software leads to better research” (Goble, 2014). Our work, thus, can provide insights to support the R developers, superficially in the context of program comprehension and updating comments, which could lead to better maintainability. This is one of the two important aspects for scientific software developers, and R is considered as a language used for the scientific software development (Arvanitou et al., 2021). We provide discussions and implication at the end, that helps shaping the future research for code intelligence in R, including the need to curate R datasets and developing separate techniques to increase the performance of the models, including LLMs, for the two R styles.
@@ 2.1. Pre-trained Language Model for Code
Pre-trained models, specifically transformer-based models (Vaswani et al., 2017), have found significant applications in various software engineering tasks including code summarization, bug detection, and code translation (Feng et al., 2020; Guo et al., 2022). Examples of these models are CodeBERT (Feng et al., 2020), GraphCodeBERT  (Guo et al., 2020), CodeT5 (Wang et al., 2021b), and UniXCoder (Guo et al., 2022). PLMs outperform many of the state-of-the-art models on several datasets and tasks and have been used extensively in multiple software engineering tasks (Feng et al., 2020; Husain et al., 2019; Niu et al., 2023).
@@
Due to the advantages of the PLMs for source code, there are works that investigate the usage and capabilities of Code-PLMs. Wan et al. (2022) studied the features captured by CodeBERT and GraphBERT through attention analysis, probing, and syntax tree induction, on Python, Java, and PHP. Troshin and Chirkova (2022) introduce probing tasks to discover the syntactic and semantic information of various Code-PLMs. Karmakar and Robbes (2021) uses probing tasks to study the Code-PLMs but does not include R language. Ahmed et al. (2023b) studies the Code-PLMs specifically by evaluating the models’ ability to learn semantics through objective and straightforward evaluation.
@@ 2.2. Scientific Programming Languages
Scientific programming languages play a pivotal role in the realm of data analysis, statistical modeling, and visualization, catering to the needs of researchers, statisticians, and data scientists (Staples, 2023). Among the diverse array of programming languages utilized in scientific domains, R stands out as a powerful and popular choice due to its robust statistical capabilities and extensive ecosystem (Giorgi et al., 2022). One of the key contributors to R’s widespread adoption is its vibrant and dedicated community, which has actively expanded the language’s functionality. Over the past decade, there has been a lot of research devoted to expanding the applicability and functionality of the R language. Notable contributions include the development of ggplot2 (Villanueva and Chen, 2019), a sophisticated and flexible data visualization package that enables the creation of high-quality graphics with ease. The Tidy-verse (Wickham et al., 2019), a collection of R packages designed to enhance data manipulation and analysis, has further streamlined workflows, promoting code readability and efficiency. RStudio, an integrated development environment (IDE) specifically tailored for R, has played a pivotal role in facilitating a user-friendly programming experience. Offering features like syntax highlighting, code completion, and interactive data visualization. Benefiting from the fact that these contributions have made the R language more applicable, many researchers from other fields have incorporated the R language into their analyses (de Jong et al., 2022; Zhao et al., 2021). Staples (2023) studied how R evolved, and showed that the emergence of the Tidy-verse is one of the primary causes behind the evolution of R. At the time of writing this manuscript, there is no research on the applicability and commonality of pre-trained models and large language models over R language.
@@ 2.3. Scientific Software, Jupyter Notebooks, and R Studies
In software engineering, several attempts and perspectives have been aimed at advancing code intelligence in R. However, to the best of our knowledge, no research has been conducted explicitly exploring the potential of pre-trained language models in this context.
Popov et al. (2021) develop a deep learning model for code completion in R. Sharma et al. (2022b) studied Self-Admitted Technical Debt (SATD) in R software packages and Chandramouli et al. (2022) presented analyzeR, a SonarQube plugin for analyzing object-oriented R packages. Zanella and Liu (2020) study the R packages on CRAN through a social network perspective. They report that the performance of R packages can be explained as a flow of information in the obtained network. Huang and Yang (2022) conducts an empirical study on the R ecosystem and found that the R development has been done by also software engineers and from other disciplines such as biology, environmental and medical science, other than statistics.
In another work, Plakidas et al. (2017) investigates the evolution of R and Ooms (2013) discusses the dependency versioning of R and demonstrates the issue and the solution through three use cases. Morandat et al. (2012) applies static and dynamic analysis to assess the design of R. Korkmaz et al. (2018) studies how the impact of R packages could be modeled through their dependencies and their contributors’ network.
@@
Another closely researched area is scientific software and scientific computing. Ivie and Thain (2018) discussed the reproducibility in scientific computing and Goble (2014) reports that 91% of scientists surveyed online use scientific software, and at least 38% spend one fifth of their time developing scientific software. The author discusses that better software leads to better research.
Other works also study the data science practices in Jupyter notebooks. Liu et al. (2021) develop a model to generate code documentation for Jupyter Notebooks. Zhang et al. (2022) collect a dataset and build a model to classify the cells of Jupyter notebooks in data science projects. Yang et al. (2021) develop a Jupyter notebook extension based on program synthesis and and test selection to generate documentations that help users find bugs for data wrangling. Dong et al. (2021) study the data cleaning practices in 30 Jupyter notebooks. Robinson et al. (2022) examine Jupyter notebooks to define error identification strategies of Python notebooks. Their work is a replication of a similar work with user study to find these strategies for R Markdown notebooks Abdel-Ghani et al. (2022). Maintenance (Jiang et al., 2022), bug analysis (de Santana et al., 2022), structure improvement (Titov et al., 2022), and understanding documentation practices Wang et al. (2021a) are among other research explored in Jupyter notebooks.
Differences among our work and the current literature. Among the related works, the closest ones are the approaches that use deep learning to detect SATD (Sharma et al., 2022b) or complete code in R (Popov et al., 2021). The work multiple_E (Cassano et al., 2022) purposed a translater that is able to translate HumanEval (Chen et al., 2021) benchmark dataset to other programming language including R. Still, none of these or other research studies the ability of Code-PLMs for different tasks in R or considers the two R styles. Though the Jupyter notebook supports R programming language, the current studies mainly focus on Python-based libraries, excluding R. The other research on R programming language are related to empirical studies to gain a better insight about its different aspects. Our work differs with all the current literature, by first, providing a dataset for code summarization and method name prediction, and then exploring the capabilities of Code-PLMs for R programming language, in various settings including intra- and cross-projects and considering its different styles, Tidy-verse and Base.
@@ 3. Methodology
In this section, we will review the research questions, data collection process, studied tasks and Code-PLMs, along with our approach to answer each of the research questions.
@@ 3.1. Research Questions
Many studies show that Code-PLMs obtain state-of-the-art performance for programming languages on the CodeSearchNet dataset. However, to what extent these models can accomplish the tasks for R language is not defined. Therefore, in this RQ, we evaluate the performance of the Code-PLMs for R. The answer to this RQ reveals the capability of the code-PLMs for R language, which is not explored in the existing research.
@@
Previous research (LeClair and McMillan, 2019; Ahmed and Devanbu, 2023) studied the effect of different dataset designs on the performance of the code summarization models. In this RQ, we investigate the outcome of models based on choosing the training and test functions from the same or different R repositories, also known as intra- or cross-project data selection. Although intra- or cross-project effect has been explored for other languages, it has not been explored for R. The results can shed light on designing experiments and the dataset to increase the performance of the models.
@@
Recent advances in foundation models or Large Language Models have shown promising results for some of the software engineering tasks (Brown et al., 2020; Roziere et al., 2023; Wang et al., 2021b) Accordingly, in this RQ, we will explore the performance of LLMs for the studied tasks for R. Currently, there is limited research on assessing the capabilities of LLMs for low-resource languages, including R. With the interest in using LLMs, the results of RQ5 will shed light on the potential of these models for code intelligence in R, compared to the code-PLMs.
@@ 3.2.1. Overview
Figure 1 presents an overview of our methodology for preparing the data. First, we collect GitHub repositories that include R files, considering the inclusion/exclusion criteria that we discuss next. Then, the eligible code snippets are parsed into code tokens using the tree-sitter parser. In the end, we followed the Roxygen2 documentation structure to match the code snippets and their corresponding natural language description to generate a dataset where each line of the data contains a pair of Programming Language (PL) and its correspondence Natural Language (NL).
@@ 3.2.2. Scraping GitHub Repositories
  • The project should be a non-personal project. So all personal projects are excluded.
  • The project should not be archived. When a repository is archived, its issues, pull requests, code, labels and other components of the project become read-only. The owner and collaborators also can not update the project. So all archived repositories are excluded.
@@ 3.2.3. Data
To create the dataset for software engineering tasks, we formatted our data following the CodeSearchNet dataset (Husain et al., 2019) structure, where each line of the JSON file contains one data example, including the function name, the code tokens, the natural language description present in their respective documentation and the doc Strings. Each data point also contains metadata on the function, such as the repository it was extracted from. In order to do so, we performed a series of preprocessing and filtering on the raw data. The following discusses preprocessing and filtering in detail.
A total of $1,664$ GitHub repositories were collected using the GitHub API following the inclusion and exclusion criteria discussed above. We then pre-processed the data to obtain the required input for the software engineering tasks. Roxygen2888https://roxygen2.r-lib.org/, a popular documentation system for R that is similar to JavaDoc, was used to parse the functions into Abstract Syntax Trees (ASTs) and comments. We first make sure that these repositories are written in R and have a Roxygen framework. The Roxygen framework is needed to ensure the code comments could be extracted. For this purpose, we first inspect the file types of the source codes in our collected data. If the package includes .R files, typically found in an R folder in the root directory, it indicates the use of R language. Then, we look for Roxygen2 comments within these .R files. Roxygen2 is widely used and it manifests in the source code as specific comments, starting with the ‘#’ symbol, followed by the ‘@’ symbol, like ‘# @param’ or ‘# @return’. If we observe such comments, it suggests that the package employs Roxygen2 for documentation. Lastly, we examine the DESCRIPTION file in the package’s root directory. This file contains various metadata about the package, such as version number, maintainer, description, and package dependencies. If Roxygen2 is listed in the Imports or Suggests field, it provides a further indication that the R package employs Roxygen2. This process allowed us to extract the necessary information from the R code to form the input for the Pre-trained-models.

### v3

@@ 1. Introduction
However, even though the R language has unparalleled importance in related fields, very limited research on code intelligence in R has been conducted. A conventional GPT-2 (Radford et al., 2018) model is used for the code completion task in R (Popov et al., 2021). Though a large number of pre-trained models have been successfully developed in recent years and have proven their effectiveness in various software engineering (SE) tasks (Liu et al., 2019; Devlin et al., 2019; Feng et al., 2020; Guo et al., 2020; Wang et al., 2021b), to the best of our knowledge, no model has attempted to extend their effectiveness to the R language. While progress has been made in code intelligence for R (Popov et al., 2021), there exists a gap in the literature, limiting the realization of its full potential.
@@
In this study, we intend to explore to which the current techniques are applicable to R. However, as there is no current dataset, we first collect a dataset of high-quality R packages from GitHub repositories (inclusion/exclusion criteria described in Section 3.2.2). Then, we consider four pre-trained models, CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2020), UniXCoder (Guo et al., 2022), and CodeT5 (Wang et al., 2021b) to empirically study their performance for two tasks of code summarization and method name prediction, with various programming styles and settings. The results of the study are compared across models and with other programming languages. We then investigate an R-specific approach based on different programming styles to evaluate its effect on the models’ performance. In this study we also investigate the performance of LLMs for the R language on code generation. This paper is the second Stage of our accepted TOSEM Registered Paper [^4] https://dl.acm.org/journal/tosem/registered-papers, in which we are reporting the results along with the discussions.
Our results reveal several key findings regarding the models and their capabilities for the R programming language. First, the models studied consistently experience varying degrees of performance degradation when processing R code compared to other programming languages. Second, our experiments demonstrate that multi-language fine-tuning does not uniformly improve performance across all PLMs for R-specific tasks, contrary to expectations. Third, we find that different syntax of the R styles in Tidy-verse and Base code significantly impact the performance, which affects the results more for the code summarization task. The challenges extend beyond syntax to project-specific contexts inherent in R codebases, which pose a substantial challenge when attempting cross-project training. This is further complicated by the observation that combined syntax approaches generally underperform compared to either individual syntax paradigms. Our experiments with LLMs also show the low Pass@K scores for code generation, falling much behind the scores obtained for Python. These insights emphasize the need for more nuanced approaches to improving model performance for R programming, beyond simply increasing training data or applying multi-language fine-tuning strategies.
@@
It is worth noting that though the foundation models (i.e., Large Language Models) in the past few months have found their way into SE, there is rarely a study or application available to use them for R. Additionally, their usage is costly if one wants to call their APIs. Therefore, their usage is not available for all, and fine-tuning them is not possible due to being large, besides their hidden ability for this programming language. LLMs have some shortcomings (Miceli-Barone et al., 2023), and still studying the smaller Code-PLMs can benefit the community. Therefore, we have studied the Code-PLMs and also added results for code generation using LLMs.
Significance: This study provides valuable insights into the effectiveness of using pre-trained code models for R and highlights the need for further research in this field. The results of our work can be used as a reference for future research and provide a foundation for the development of intelligent tools for the R community. This is of specific importance as research shows that many scientists develop/use scientific software and “better software leads to better research” (Goble, 2014). Our work thus can provide insights to support R developers, superficially in the context of program comprehension and updating comments, which could lead to better maintainability. This is one of the two important aspects for scientific software developers, and R is considered as a language used for scientific software development (Arvanitou et al., 2021). We provide discussions and implications at the end, which help shape the future research for code intelligence in R, including the need to curate R datasets and develop separate techniques to increase the performance of the models, including LLMs, for the two R styles.
@@ 2.1. Pre-trained Language Model for Code
Pre-trained models, specifically transformer-based models (Vaswani et al., 2017), have found significant applications in various software engineering tasks, including code summarization, bug detection, and code translation (Feng et al., 2020; Guo et al., 2022). Examples of these models are CodeBERT (Feng et al., 2020), GraphCodeBERT  (Guo et al., 2020), CodeT5 (Wang et al., 2021b), and UniXCoder (Guo et al., 2022). PLMs outperform many of the state-of-the-art models on several datasets and tasks and have been used extensively in multiple software engineering tasks (Feng et al., 2020; Husain et al., 2019; Niu et al., 2023).
@@
Due to the advantages of the PLMs for source code, there are works that investigate the usage and capabilities of Code-PLMs. Wan et al. (2022) studied the features captured by CodeBERT and GraphBERT through attention analysis, probing, and syntax tree induction, on Python, Java, and PHP. Troshin and Chirkova (2022) introduces probing tasks to discover the syntactic and semantic information of various Code-PLMs. Karmakar and Robbes (2021) uses probing tasks to study the Code-PLMs, but does not include the R language. Ahmed et al. (2023b) studies the Code-PLMs specifically by evaluating the models’ ability to learn semantics through objective and straightforward evaluation.
@@ 2.2. Scientific Programming Languages
Scientific programming languages play a pivotal role in the realm of data analysis, statistical modeling, and visualization, catering to the needs of researchers, statisticians, and data scientists (Staples, 2023). Among the diverse array of programming languages utilized in scientific domains, R stands out as a powerful and popular choice due to its robust statistical capabilities and extensive ecosystem (Giorgi et al., 2022). One of the key contributors to R’s widespread adoption is its vibrant and dedicated community, which has actively expanded the language’s functionality. Over the past decade, there has been a lot of research devoted to expanding the applicability and functionality of the R language. Notable contributions include the development of ggplot2 (Villanueva and Chen, 2019), a sophisticated and flexible data visualization package that enables the creation of high-quality graphics with ease. The Tidy-verse (Wickham et al., 2019), a collection of R packages designed to enhance data manipulation and analysis, has further streamlined workflows, promoting code readability and efficiency. RStudio, an integrated development environment (IDE) specifically tailored for R, has played a pivotal role in facilitating a user-friendly programming experience. Offering features like syntax highlighting, code completion, and interactive data visualization. Benefiting from the fact that these contributions have made the R language more applicable, many researchers from other fields have incorporated the R language into their analyses (de Jong et al., 2022; Zhao et al., 2021). Staples (2023) studied how R evolved, and showed that the emergence of the Tidy-verse is one of the primary causes behind the evolution of R. At the time of writing this manuscript, there is no research on the applicability and commonality of pre-trained models and large language models over the R language.
@@ 2.3. Scientific Software, Jupyter Notebooks, and R Studies
Recent research has shown substantial progress in code generation for popular scientific programming languages like Python and Julia. Due to Python’s extensive open-source ecosystem, large language models like Codex (Chen et al., 2021) have leveraged a large number of repositories to train and achieve high performance on widely adopted benchmarks (e.g., HumanEval (Chen et al., 2021) and MultiPL-E (Cassano et al., 2023)). Similarly, Julia benefits from growing popularity in high-performance computing and data science, making it a recurring target in the state-of-the-art multilingual code-generation evaluations (Athiwaratkun et al., 2023).
The challenges of R in this landscape are derived from its design and the constraints of the ecosystem. Although functional meta-programming via packages like rlang  [^6] https://rlang.r-lib.org/index.html offers flexibility, the R language presents unique challenges. R is a mainstay for statistical computing and data analysis, but R’s public codebase is comparatively smaller, and many open-source R repositories cater to highly specialized use cases. As a result, LLMs trained on general-purpose code datasets may struggle to capture R-specific paradigms (e.g., vectorized operations, specialized data-wrangling functions), leading to lower performance relative to Python and Julia (Chen et al., 2021). This discrepancy highlights the problem of “low-resource languages” in code generation, where the scarcity of high-quality training data hinders a model’s performance (Cassano et al., 2024). In software engineering, several attempts and perspectives have been aimed at advancing code intelligence in R. However, to the best of our knowledge, no research has been conducted explicitly exploring the potential of pre-trained language models in this context.
Popov et al. (2021) develop a deep learning model for code completion in R. Sharma et al. (2022b) studied Self-Admitted Technical Debt (SATD) in R software packages, and Chandramouli et al. (2022) presented analyzeR, a SonarQube plugin for analyzing object-oriented R packages. Zanella and Liu (2020) study the R packages on CRAN through a social network perspective. They report that the performance of R packages can be explained as a flow of information in the obtained network. Huang and Yang (2022) conducts an empirical study on the R ecosystem and found that the R development has been done by software engineers and from other disciplines such as biology, environmental, and medical science, other than statistics.
In another work, Plakidas et al. (2017) investigates the evolution of R, and Ooms (2013) discusses the dependency versioning of R and demonstrates the issue and the solution through three use cases. Morandat et al. (2012) applies static and dynamic analysis to assess the design of R. Korkmaz et al. (2018) studies how the impact of R packages could be modeled through their dependencies and their contributors’ network.
@@
Another closely researched area is scientific software and scientific computing. Ivie and Thain (2018) discussed the reproducibility in scientific computing and Goble (2014) reports that 91% of scientists surveyed online use scientific software, and at least 38% spend one-fifth of their time developing scientific software. The author discusses that better software leads to better research.
Other works also study the data science practices in Jupyter notebooks. Liu et al. (2021) develops a model to generate code documentation for Jupyter Notebooks. Zhang et al. (2022) collects a dataset and builds a model to classify the cells of Jupyter notebooks in data science projects. Yang et al. (2021) develop a Jupyter notebook extension based on program synthesis and and test selection to generate documentations that help users find bugs for data wrangling. Dong et al. (2021) study the data cleaning practices in 30 Jupyter notebooks. Robinson et al. (2022) examine Jupyter notebooks to define error identification strategies of Python notebooks. Their work is a replication of a similar work with a user study to find these strategies for R Markdown notebooks Abdel-Ghani et al. (2022). Maintenance (Jiang et al., 2022), bug analysis (de Santana et al., 2022), structure improvement (Titov et al., 2022), and understanding documentation practices Wang et al. (2021a) are among other research explored in Jupyter notebooks.
Differences among our work and the current literature. Among the related works, the closest ones are the approaches that use deep learning to detect SATD (Sharma et al., 2022b) or complete code in R (Popov et al., 2021). The work multiple_E (Cassano et al., 2022) proposed a translator that is able to translate the HumanEval (Chen et al., 2021) benchmark dataset to other programming languages, including R. Still, none of these or other research studies the ability of Code-PLMs for different tasks in R or considers the two R styles. Though the Jupyter notebook supports the R programming language, the current studies mainly focus on Python-based libraries, excluding R. The other research on the R programming language is related to empirical studies to gain a better insight into its different aspects. Our work differs from all the current literature by first providing a dataset for code summarization and method name prediction, and then exploring the capabilities of Code-PLMs for the R programming language, in various settings including intra- and cross-projects and considering its different styles, Tidy-verse and Base.
@@ 3. Methodology
In this section, we will review the research questions, data collection process, studied tasks, and Code-PLMs, along with our approach to answer each of the research questions.
@@ 3.1. Research Questions
Many studies show that Code-PLMs obtain state-of-the-art performance for programming languages on the CodeSearchNet dataset. However, to what extent these models can accomplish the tasks for the R language is not defined. Therefore, in this RQ, we evaluate the performance of the Code-PLMs for R. The answer to this RQ reveals the capability of the code-PLMs for the R language, which is not explored in the existing research.
@@
Previous research (LeClair and McMillan, 2019; Ahmed and Devanbu, 2023) studied the effect of different dataset designs on the performance of the code summarization models. In this RQ, we investigate the outcome of models based on choosing the training and test functions from the same or different R repositories, also known as intra- or cross-project data selection. Although intra- or cross-project effects have been explored for other languages, it has not been explored for R. The results can shed light on designing experiments and the dataset to increase the performance of the models.
@@
Recent advances in foundation models or Large Language Models have shown promising results for some of the software engineering tasks (Brown et al., 2020; Roziere et al., 2023; Wang et al., 2021b). Accordingly, in this RQ, we will explore the performance of LLMs for the studied tasks for R. Currently, there is limited research on assessing the capabilities of LLMs for low-resource languages, including R. With the interest in using LLMs, the results of RQ5 will shed light on the potential of these models for code intelligence in R, compared to the code-PLMs.
@@ 3.2.1. Overview
Figure 1 presents an overview of our methodology for preparing the data. First, we collect GitHub repositories that include R files, considering the inclusion/exclusion criteria that we discuss next. Then, the eligible code snippets are parsed into code tokens using the tree-sitter parser. In the end, we followed the Roxygen2 documentation structure to match the code snippets and their corresponding natural language description to generate a dataset where each line of the data contains a pair of Programming Language (PL) and its corresponding Natural Language (NL).
@@ 3.2.2. Scraping GitHub Repositories
  • The project should be a non-personal project. So, all personal projects are excluded.
  • The project should not be archived. When a repository is archived, its issues, pull requests, code, labels, and other components of the project become read-only. The owner and collaborators also can not update the project. Therefore, all archived repositories are excluded.
@@ 3.2.3. Data
To create the dataset for software engineering tasks, we formatted our data following the CodeSearchNet dataset (Husain et al., 2019) structure, where each line of the JSON file contains one data example, including the function name, the code tokens, the natural language description present in their respective documentation, and the doc Strings. Each data point also contains metadata on the function, such as the repository it was extracted from. In order to do so, we performed a series of preprocessing and filtering on the raw data. The following discusses preprocessing and filtering in detail.
A total of $1,664$ GitHub repositories were collected using the GitHub API following the inclusion and exclusion criteria discussed above. We then pre-processed the data to obtain the required input for the software engineering tasks. Roxygen2999https://roxygen2.r-lib.org/, a popular documentation system for R that is similar to JavaDoc, was used to parse the functions into Abstract Syntax Trees (ASTs) and comments. We first make sure that these repositories are written in R and have an Roxygen framework. The Roxygen framework is needed to ensure that the code comments can be extracted. For this purpose, we first inspect the file types of the source codes in our collected data. If the package includes .R files, typically found in an R folder in the root directory, it indicates the use of the R language. Then, we look for Roxygen2 comments within these .R files. Roxygen2 is widely used and it manifests in the source code as specific comments, starting with the ‘#’ symbol, followed by the ‘@’ symbol, like ‘# @param’ or ‘# @return’. If we observe such comments, it suggests that the package employs Roxygen2 for documentation. Lastly, we examine the DESCRIPTION file in the package’s root directory. This file contains various metadata about the package, such as version number, maintainer, description, and package dependencies. If Roxygen2 is listed in the Imports or Suggests field, it provides a further indication that the R package employs Roxygen2. This process allowed us to extract the necessary information from the R code to form the input for the pre-trained models.